Use the speed of light as c = 3 × 10^8 m/s when needed.
3. It takes light 1.25 seconds to reach the moon from the earth. How far away is the Moon (in meters)?

Answer by Alan3354(69443)

Use the speed of light as c = 3 × 10^8 m/s when needed.
3. It takes light 1.25 seconds to reach the moon from the earth. How far away is the Moon (in meters)?
-----------
d = r*t
d = 3*10^8*1.25 meters